# Designing and Contextualising Probes for African Languages

### v1

@@ 3.3 NTC and NER Probe Baselines
Control tasks define word-level control behaviour, so they are not applicable to sentence-level tasks. To interpret NTC probes, we compare the performance of probes trained on PLMs to those trained on random contextual representations. Following Hewitt and Manning (2019), we use an untrained bi-LSTM, mean-pooling word-level outputs to produce a single sentence representation. Probes trained on these outputs can leverage contextual information, but no linguistic knowledge.
@@ 4.2 Languages
The six language in our study (Swahili, Igbo, Hausa, Luganda, isiXhosa, and Naija Pidgin) are included in all three Masakhane [^1] https://www.masakhane.io/ datasets. We chose these languages specifically to cover several language families, a broad range of data availability, and varying levels of representation in existing PLMs (as shown in Table 1). As shown in Table 2, the our languages covers four language families across East, West, and Southern Africa.
@@ 4.3 PLMs
The seven PLMs select for our study represent established approaches to developing PLMs for African languages. XLM-R-base and XLM-R-large (Conneau et al., 2020) employ massively multilingual pretraining, while AfroXLMR-base, AfroXLMR-large (Alabi et al., 2022), and Nguni-XLMR-large (Meyer et al., 2024) adapt XLM-R to a more limited set of African languages. AfroXLMR takes a broader adaptation approach than Nguni-XLMR, which focusses only on the four Nguni languages, a group of related languages which includes isiXhosa. AfriBERTa (Ogueji et al., 2021) represents memory-efficient pretraining – it is our smallest model both in terms of parameters and training data size. AfroLM (Dossou et al., 2022) represents sample-efficient pretraining, since it’s self-active learning maximises available data by identifying beneficial training samples.
@@ 5.4 Analysing Trends Across Tasks
Multilingual adaptation is known to be a reliable method to improve downstream performance for low-resource languages. Our results show that this being achieved, in part, by enhancing internal representations of target languages. AfroXLMR-large and Nguni-XLMR-large have the most instances of top-performing layers (six cases each). AfroXLMR-large performs best across Swahili, Igbo, and Hausa, all three of which belong to different language families. The multilingual adaptation of AfroXLMR is highly effective at enhancing linguistic feature encoding across typologically diverse languages. Nguni-XLMR-large performs best for isiXhosa and Luganda, which is also of the Bantu language family. The more focussed, linguistically oriented adaptation of Nguni-XLMR effectively enhances linguistic knowledge for a more limited set of related languages.

### v2

@@ 3.3 NTC and NER Probe Baselines
Control tasks define word-level control behaviour, so they are not applicable to sentence-level tasks. To interpret NTC probes, we compare the performance of probes trained on PLMs to those trained on random contextual representations. Following Hewitt and Manning (2019), we use an untrained bi-LSTM, and mean-pooling word-level outputs to produce a single sentence representation. Probes trained on these outputs can leverage contextual information, but no linguistic knowledge.
@@ 4.2 Languages
The six language in our study (Swahili, Igbo, Hausa, Luganda, isiXhosa, and Naija Pidgin) are included in all three Masakhane [^1] https://www.masakhane.io/ datasets. We chose these languages specifically to cover several language families, a broad range of data availability, and varying levels of representation in existing PLMs (as shown in Table 1). As shown in Table 2, the languages cover four language families across East, West, and Southern Africa.
@@ 4.3 PLMs
The seven PLMs selected for our study represent established approaches to developing PLMs for African languages. XLM-R-base and XLM-R-large (Conneau et al., 2020) employ massively multilingual pretraining, while AfroXLMR-base, AfroXLMR-large (Alabi et al., 2022), and Nguni-XLMR-large (Meyer et al., 2024) adapt XLM-R to a more limited set of African languages. AfroXLMR takes a broader adaptation approach than Nguni-XLMR, which focusses only on the four Nguni languages, a group of related languages which includes isiXhosa. AfriBERTa (Ogueji et al., 2021) represents memory-efficient pretraining – it is our smallest model both in terms of parameters and training data size. AfroLM (Dossou et al., 2022) represents sample-efficient pretraining, since its self-active learning maximizes available data by identifying beneficial training samples.
@@ 5.4 Analysing Trends Across Tasks
Multilingual adaptation is known to be a reliable method to improve downstream performance for low-resource languages. Our results show that this is being achieved, in part, by enhancing internal representations of target languages. AfroXLMR-large and Nguni-XLMR-large have the most instances of top-performing layers (six cases each). AfroXLMR-large performs best across Swahili, Igbo, and Hausa, all three of which belong to different language families. The multilingual adaptation of AfroXLMR is highly effective at enhancing linguistic feature encoding across typologically diverse languages. Nguni-XLMR-large performs best for isiXhosa and Luganda (which is also of the Bantu language family). The more focussed, linguistically oriented adaptation of Nguni-XLMR effectively enhances linguistic knowledge for a more limited set of related languages.